OpenStudy (anonymous):

Create an equation. Use the graph below to create the equation of the rainbow parabola.

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

so wht do u thiink wht r the roots of the equation???

OpenStudy (anonymous):

um -6 and 6

rishavraj (rishavraj):

cool......xD means u don't need any help :))

OpenStudy (anonymous):

lol! I so I got -x^2+36 ??

OpenStudy (jdoe0001):

well... notice whre the vertex is at the vertex is at (0, 36) well.. actaully that's quite close, one sec

OpenStudy (anonymous):

and there's a 2nd part that I'm not understading

OpenStudy (anonymous):

Create a table of values for a linear function. A drone is in the distance, flying upward in a straight line. It intersects the rainbow at two points. Choose the points where your drone intersects the parabola and create a table of at least four values for the function. Remember to include the two points of intersection in your table.

OpenStudy (jdoe0001):

\(\bf y=a(x-{\color{brown}{ h}})^2+{\color{blue}{ k}} \qquad vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\) the parabola is upside-down, meaning the "a" is negative, yes thus \(\bf y=a(x-{\color{brown}{ h}})^2+{\color{blue}{ k}} \qquad vertex\ ({\color{brown}{ h}},{\color{blue}{ k}}) \\ \quad \\ y=-a(x-0)^2+36\implies y=-ax^2+36\) we know the roots, thus, when y =0 , x = 6 or -6, let's use 6 so 6, 0 then \(\bf y=-ax^2+36\implies \cfrac{y-36}{-x^2}=a\implies \cfrac{0-36}{-6^2}=a\)

OpenStudy (anonymous):

Yes, so it would be -x^2+36 :)

OpenStudy (jdoe0001):

ohh yeah, it would, anyhow, kudos :)

OpenStudy (jdoe0001):

lemme check the other bit

OpenStudy (anonymous):

Thanks:) Because I have no idea what to do

OpenStudy (jdoe0001):

well, is a system of equations lemme see if hmmm lemme graph it

OpenStudy (anonymous):

Ok.

OpenStudy (jdoe0001):

http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiIteF4yKzM2IiwiY29sb3IiOiIjRTMxMjEyIn0seyJ0eXBlIjowLCJlcSI6IigxLzMpeCsyNSIsImNvbG9yIjoiIzIyMUJFMCJ9LHsidHlwZSI6MTAwMCwid2luZG93IjpbIi0xMy40MjQ4MzUyMDUwNzgxMTgiLCIxMS45NjU3ODk3OTQ5MjE4NjgiLCIyMS4xMDg2MjczMTkzMzU5MzgiLCIzNi43MzM2MjczMTkzMzU5MyJdfV0- you have the "rainbow" parabola and then you're asked to make a LINEar function you're expected to use a table of values for the linear one and to include at least, four values, more is fine so if you use 10 values is ok as well, but not less than 4 values for the table notice the graph.... the LINEar equation graph, crosses the parabola at two points

OpenStudy (anonymous):

how did you get 1/3x+25?

OpenStudy (jdoe0001):

made up :), you're expected to make one up, thus

OpenStudy (jdoe0001):

I simply made sure the line was secant to the parabola

OpenStudy (jdoe0001):

but you can cross the parabola by other lines as well you could simply pick two points, to make a line one on the left, one on the right make sure they "cut" the parabola through, so they touch it at two points and with two points you can get the equation of that line from there you can get more values for the table

OpenStudy (anonymous):

ok, so what would the points be?

OpenStudy (jdoe0001):

well.. let's use the 1/3x + 25 one sec

OpenStudy (anonymous):

why couldn't we use like 9/5x+20 or 2x+27? And okay, ill wait:)

OpenStudy (jdoe0001):

either one will work, so long they "cut through" the parabola

OpenStudy (anonymous):

okay:) So 9/5x+20 would cut through right?

OpenStudy (anonymous):

Thank you:)! And to find the points, we could just plug in numbers? or does it have to be specific

OpenStudy (jdoe0001):

\(\begin{array}{rrllll} x&y\\ \\\hline\\ -6&\frac{1}{3}\cdot -6+36\\ -4&\frac{1}{3}\cdot -4+36\\ -2&\frac{1}{3}\cdot -2+36\\ 0&\frac{1}{3}\cdot 0+36\\ 2&\frac{1}{3}\cdot 2+36\\ 4&\frac{1}{3}\cdot 4+36\\ 6&\frac{1}{3}\cdot 6+36\\ \end{array}\) there, that'd a table for the blue one

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Yayyy! :) Lol I finally got it:) Thanks so muchhh

OpenStudy (jdoe0001):

np